OpenStudy (anonymous):

how do you put this into simplified radical form? the square root of 25tu^11?

OpenStudy (mertsj):

I don't see a radical anywhere

OpenStudy (anonymous):

woops...it should the square root of all that..sorry

OpenStudy (mertsj):

\[\sqrt{32x^7y^6}=\sqrt{16(2)x^6xy^6}=4x^3y^3\sqrt{2x}\]

OpenStudy (anonymous):

what about the square root of 25tu^11?

OpenStudy (mertsj):

\[5u^5\sqrt{tu}\]